# Do spectral cues matter in contrast-based graph self-supervised learning?

### v1

@@ 5 Edge perturbation is all you need
So far, our findings indicate that spectral augmentation is not particularly effective in contrast-based graph self-supervised learning. It may suggest that spectral augmentation essentially amounts to random topology perturbation, based on significant contradictions in previous studies [16, 18, 36] and the theoretical insight that a shallow encoder can hardly capture spectral properties. In fact, most of the spectral augmentations are basically performing edge perturbations on the graph with some targeted directions. Since we preliminarily conclude that it is quite difficult for those augmentations to benefit from the spectral properties of graphs, it is very intuitive to hypothesize that edge perturbation itself matters in the learning process.
Consequently, we are turning back to Edge Perturbation (EP), a more straightforward and proven method for augmenting graph data. The two primary methods of edge perturbation are DropEdge and AddEdge. We want to claim that edge perturbation has a better performance than spectral augmentations and prove empirically that none of them actually or even can benefit much from any spectral information and properties. Also, we demonstrate edge perturbation is much more efficient in practical applications both for time and space sake, where spectral operations are almost infeasible. Overall, we will support the idea with evidence in the following sections that simple edge perturbation is not only good enough but even very optimal in contrast-based graph SSL and we don’t need to care about much for spectral cues.
@@
DropEdge. Edge dropping is the process of randomly removing a subset of edges from the original graph to create an augmented view. Adopting the definition from [26], let $\mathcal{G}=(\mathbf{A},\mathbf{X})$ be the original graph with adjacency matrix $\mathbf{A}$. We introduce a mask matrix $\mathbf{M}$ of the same dimensions as $\mathbf{A}$, where each entry $M_{ij}$ follows a Bernoulli distribution with parameter $1-p$ (denoted as the drop rate). The edge-dropped graph $G^{\prime}$ is then obtained by element-wise multiplication of $\mathbf{A}$ with $\mathbf{M}$:
@@
where $\odot$ denotes the Hadamard product (element-wise multiplication).
@@ 5.1 Advantage of edge perturbation over spectral augmentations
Edge perturbation offers several key advantages over spectral augmentation, making it a more effective and practical choice for contrast-based graph self-supervised learning. Compared to augmentations related to the graph spectrum, it has three major advantages.
@@
Significantly better efficiency. Edge perturbation methods such as edge dropping (DropEdge) and edge adding (AddEdge) are computationally efficient. Unlike spectral augmentation, which requires costly eigenvalue and eigenvector computations, edge perturbation can be implemented with basic graph operations. This efficiency translates to faster training and inference times, making it more suitable for large-scale graph datasets and real-time applications. As shown in Table 1, the time and space complexity of spectrum-related calculations are several orders of magnitude higher than those of simple edge perturbation operations. This makes spectrum-related calculations impractical for the large datasets typically encountered in real-world applications.
@@
Most importantly and directly, our comprehensive empirical studies indicate that edge perturbation methods lead to significant improvements in model performance, as presented and analyzed in Sec. 6. From the results there, the conclusion can be drawn that the performance of the proposed augmentations is not only better than those of spectrum-related competitors but also matches or even surpasses the performance of other strong benchmarks.
@@ 6 Experiments on SSL performance
6.1 Experimental Settings.
Task and Datasets. We conducted extensive experiments for node-level classification on seven datasets: Cora, CiteSeer, PubMed [13], Photo, Computers [27], Coauthor-CS, and Coauthor-Phy. These datasets include various types of graphs, such as citation networks, co-purchase networks, and co-authored networks. Additionally, we carried out graph-level classification on five datasets from the TUDataset collection [21], which include biochemical molecules and social networks. More details of these datasets be found in Appendix B.
Baselines. We conducted experiments under four CG-SSL frameworks: MVGRL, GRACE, G-BT, and BGRL, using DropEdge, AddEdge, and SPAN [16] as augmentation strategies. For MVGRL, we also compared its original PPR augmentation. For the node classification task, we use GCA [43], GMI [24], DGI [30], and SpCo [18] as baselines. For the graph classification task, we use RGCL [15] and GraphCL [37] as baselines. The experimental details about the training configurations can be found in Appendix B
@@ 6.2 Experimental results
We present the results of the node classification and graph classification tasks in Table 2 and Table 3, respectively. Our comparative analysis of graph augmentation for both node and graph classification reveals distinct performance trends. For node classification, DropEdge consistently achieves the best performance across multiple datasets and CG-SSL frameworks, demonstrating superior robustness and consistency. While AddEdge also achieves competitive accuracy, DropEdge stands out in this area. In graph classification, AddEdge frequently achieves the best performance across multiple datasets and CG-SSL frameworks, showing superior and more consistent results. In contrast, SPAN generally underperforms relative to both DropEdge and AddEdge  encountering scalability issues on larger datasets and suffering from a high overhead of training time.
Table 2: Node classification. Results of baselines with $\dagger$’ are adopted directly from previous works. MVGRL+PPR is the original setting of MVGRL. The best results in each cell are highlighted by grey. The best results overall are highlighted with bold and underline.
@@
Table 3: Graph classification. Results of baselines with $\dagger$’ are adopted directly from previous works. MVGRL+PPR is the original setting of MVGRL. The best results in each cell are highlighted by grey. The best results overall are highlighted with bold and underline.
@@ 7 The insignificance of Spectral Cues
Given the superior empirical performance of edge perturbations mentioned in Sec. 6, one may still argue whether it is a result of some spectral cues or not, as all the analyses mentioned are not direct evidence of the insignificance of the spectral cues in the study. To clarify this, we have three questions to answer, (1) Can GNN encoders learn spectral information from augmented graphs produced edge perturbations? (2) Are spectrum in spectral augmentation necessary? (3) Is spectral information a significant factor in the performance of edge perturbation? Given the questions, we conduct a series of experimental studies to answer them respectively in Sec. 7.1, 7.2 and Appendix E.2.
7.1 Degeneration of the spectrum after EP
@@
Here we want to conduct studies to answer the question of whether the GNN encoders applied can learn spectral information from the augmented graph views produced by EP. Therefore, we collect the spectrum of all augmented graphs ever produced along the way of the contrastive learning process of the best framework with the optimal parameter we have in this study, i.e. G-BT + EP with best drop rate $p$ or add rate $q$, and calculate the average one for each representative dataset in this study for both node- and graph-level tasks. We find that though the average spectrum of those original graphs is strikingly different, that of augmented graphs is quite similar for node- and graph-level tasks, respectively. This indicates a certain degree of degeneration of the spectra as they are no longer easy to separate after EP. Therefore, GNN encoders can hardly learn spectral information and properties between different original graphs from those augmented graph views. Note that, though we have defined some context of frameworks, this result is generally only dependent on the augmentation methods. Due to the limited space, we will elaborate and visualize the node-level results in this section and postpone the graph-level ones in Appendix E.1, as they support the claim very consistently.
Node-Level analysis.
@@
Here, we visualize the distributions of the average spectrum of graphs at the node level using histograms. The spectral distribution for each graph is represented by a sorted vector of its eigenvalues. When referring to the average spectrum, we mean the average over the eigenvalue vectors of each augmented graph. We plot the histograms of different spectra, normalized to show the probability density. Note that eigenvalues are constrained within the range [0, 2], as we adopted the commonly used symmetrical normalization. We analyze the spectral distributions of three node classification datasets: Cora, CiteSeer, and Computers. We compare the average spectral properties of both original and augmented graphs. The augmentation method used is DropEdge, applied with optimal parameters identified for the G-BT method.
The results of the visualization are presented in Fig. 2. By comparing the spectrum distributions of original graphs for the datasets in Fig. 2(a), we can easily distinguish the spectra of the three datasets. This contrasts with the highly overlapped average spectra of all the datasets, indicating the degeneration mentioned. To support this claim, we also present the comparison of the spectra of original and augmented graphs on all three datasets in Fig. 2(c), 2(d), and 2(e), respectively, to show the obvious changes after the edge perturbations.
@@ 7.2 Spectral Perturbation
Spectral perturbation on spectral augmentation baselines.
SPAN, being a state-of-the-art spectral augmentation algorithm, demonstrated the correlation between graph spectra and model performance through designated perturbation on spectral priors. However, the effectiveness of simple edge perturbation motivated us to further investigate whether such a relationship is causational.
Specifically, for each pair of SPAN augmented graphs $\mathcal{G}^{1},\mathcal{G}^{2}$, we further augment them into $\mathcal{G}^{1}_{SPA},\mathcal{G}^{2}_{SPA}$ with our proposed SPA augmentor. The SPA-augmented training is performed under the same setup as SPAN, with graphs being SPA-augmented graphs $\mathcal{G}_{SPA}$. Experiment results in Fig 3 show that the effectiveness of graph augmentation can be preserved and, in some cases improved, even if the spectral information is destroyed.
@@
$s_{SPAN}\sim p_{\theta}\left(\boldsymbol{S}_{SPAN}\left.\right|\boldsymbol{\mathcal{G}}_{0}\right),\qquad\mathcal{G}_{SPAN}\sim p_{\phi}\left(\boldsymbol{\mathcal{G}}_{SPAN}\left.\right|\boldsymbol{S}_{SPAN}\right)$ (5)
@@
$d_{SPA}\rightarrow\infty\implies p_{\theta}\left(s_{SPA}\left.\right|\boldsymbol{\mathcal{G}}_{0}\right)\rightarrow 0\implies p_{\theta,\phi}\left(\mathcal{G}_{SPA}\left.\right|\boldsymbol{\mathcal{G}}_{0}\right)\rightarrow 0$ (6)
Although the constrained generative process in Eq. 5 does indicate some extent of causality between spectral distribution $\boldsymbol{S}$ and the spectral-augmented graph distribution $\boldsymbol{\mathcal{G}}_{SPAN}$, our experiment challenges a more essential and fundamental aspect of such reasoning: such causality exists upon pre-defined generative processes, which does not intrinsically exist in the graph distributions. Even worse, such constrained generative process is incapable of modeling the full distribution of $\boldsymbol{\mathcal{G}}_{SPAN}$ itself. In our experiment setup, all $\mathcal{G}_{SPA}$ serve as strong counter examples.
@@ 8 Conclusion
In this study, we investigate the effectiveness of spectral augmentation in contrast-based graph self-supervised learning (CG-SSL) frameworks to answer the question: Do spectral cues matter in CG-SSL? Our findings indicate that spectral information does not significantly enhance learning efficacy. Instead, simpler edge perturbation techniques, such as random edge dropping for node-level tasks and random edge adding for graph-level tasks, not only compete well but often outperform spectral augmentations. To be specific, we demonstrate that the benefits of spectral augmentation diminish with shallower networks, and edge perturbations yield superior performance in both node- and graph-level classification tasks. Additionally, GNN encoders struggle to learn spectral information from augmented graphs, and perturbing edges to alter spectral characteristics does not degrade model performance. These results challenge the current emphasis on spectral augmentation, advocating for more straightforward and effective edge perturbation techniques in CG-SSL, potentially reshaping the understanding and implementation of graph self-supervised learning.
Appendix A Limitations of the study
While our study provides compelling evidence against the necessity of spectral augmentation in contrast-based graph self-supervised learning (CG-SSL), it is important to acknowledge potential limitations that open avenues for further exploration.
  1. 1.
    Our analysis primarily focuses on specific types of edge perturbation techniques, i.e. DropEdge and AddEdge. Future research could explore a broader range of augmentation methods to validate our conclusions across different contexts.
  2. 2.
    Although we cover various state-of-the-art frameworks and tasks, the generalizability of our findings could be further strengthened by applying our approach to a wider range of datasets and frameworks.
  3. 3.
    Our conclusions are partially based on observations related to network depth. Though already adopted the most commonly used configuration among all related studies, further investigation is needed to understand how different network architectures and configurations might influence the effectiveness of spectral and edge-based augmentations.
Addressing these limitations can provide deeper insights and help refine graph self-supervised learning methodologies.
Appendix B Dataset and training configuration

### v2

@@ 5 Edge perturbation is all you need
So far, our findings indicate that spectral augmentation is not particularly effective in contrast-based graph self-supervised learning. It may suggest that spectral augmentation essentially amounts to random topology perturbation, based on inconsistencies
in previous studies [19, 21, 42] and the theoretical insight that a shallow encoder can hardly capture spectral properties. In fact, most of the spectral augmentations are basically performing edge perturbations on the graph with some targeted directions. Since we preliminarily conclude that it is quite difficult for those augmentations to benefit from the spectral properties of graphs, it is very intuitive to hypothesize that edge perturbation itself matters in the learning process.
Consequently, we are turning back to Edge Perturbation (EP), a more straightforward and proven method for augmenting graph data. The two primary methods of edge perturbation are DropEdge and AddEdge. We want to claim that edge perturbation has a better performance than spectral augmentations and prove empirically that none of them actually or even can benefit much from any spectral information and properties. Also, we demonstrate edge perturbation is much more efficient in practical applications for both time and space sake, where spectral operations are almost infeasible. Overall, we will support the idea with evidence in the following sections that simple edge perturbation is not only good enough but even very optimal in CG-SSL compared to spectral augmentations.
@@
DropEdge. Edge dropping is the process of randomly removing a subset of edges from the original graph to create an augmented view. Adopting the definition from [31], let $\mathcal{G}=(\mathbf{A},\mathbf{X})$ be the original graph with adjacency matrix $\mathbf{A}$. We introduce a mask matrix $\mathbf{M}$ of the same dimensions as $\mathbf{A}$, where each entry $M_{ij}$ follows a Bernoulli distribution with parameter $1-p$ (denoted as the drop rate). The edge-dropped graph $G^{\prime}$ is then obtained by element-wise multiplication of $\mathbf{A}$ with $\mathbf{M}$ (where $\odot$ denotes the Hadamard product):
@@ 5.1 Advantage of edge perturbation over spectral augmentations
Edge perturbation offers several key advantages over spectral augmentation, making it a more effective and practical choice for CG-SSL. Compared to spectrum-related augmentations, it has three major advantages.
@@
Significantly better efficiency. Edge perturbation methods such as edge dropping (DropEdge) and edge adding (AddEdge) are computationally efficient. Unlike spectral augmentation, which requires costly eigenvalue and eigenvector computations, edge perturbation can be implemented with basic graph operations. This efficiency translates to faster training and inference times, making it more suitable for large-scale graph datasets and real-time applications. As shown in Table 1, the time and space complexity of spectrum-related calculations are several orders of magnitude higher than those of simple edge perturbation operations. This makes spectrum-related calculations impractical for large datasets typically encountered in real-world applications.
@@
Most importantly and directly, our comprehensive empirical studies indicate that edge perturbation methods lead to significant improvements in model performance, as presented and analyzed in Sec. 6. From the results there, the conclusion can be drawn that the performance of the proposed augmentations is not only better than those of spectral augmentations but also matches or even surpasses the performance of other strong benchmarks.
@@ 6 Experiments on SSL performance
6.1 Experimental Settings
Task and Datasets. We conducted extensive experiments for node-level classification on seven datasets: Cora, CiteSeer, PubMed [16], Photo, Computers [32], Coauthor-CS, and Coauthor-Phy. These datasets include various types of graphs, such as citation networks, co-purchase networks, and co-authored networks. Note that we do not include huge-scale datasets like OGBN [14] for the high complexity of spectral augmentations. While both DropEdge and AddEdge have linear complexity that can easily run on those huge datasets, no spectral augmentation can scale to them. Additionally, we carried out graph-level classification on five datasets from the TUDataset collection [24], which include biochemical molecules and social networks. More details of these datasets be found in Appendix A.
Baselines. We conducted experiments under four CG-SSL frameworks: MVGRL, GRACE, G-BT, and BGRL (mentioned in Sec 3), using DropEdge, AddEdge, and SPAN [19] as augmentation strategies. Note that there are only three very relevant studies on spectral augmentation strategies of CG-SSL to the authors’ best knowledge, i.e., SPAN, SpCo [21] and GASSER [42]. Among them, GASSER does not have open-sourced code so we are not able to reproduce any related results, but we try our best to directly adopt the best performance reported in that study to ensure comparison is possible. Also, SpCo is only applicable to node-level tasks and its implementation is not robust enough to generalize to all the node-level datasets and CG-SSL frameworks. Therefore, we manage to include the results of all the settings that it is feasible to do, which is its original setting and the combination of GRACE and it. Given the infeasibility and inaccessibility of the two, we selected SPAN as a major baseline because it is robust and general enough to all the datasets and experimental settings while allowing the modular plug-and-play integration of edge perturbation methods, enabling a very direct angle to evaluate the effectiveness of the spectral augmentations compared to much simpler alternatives.
Besides the major baselines mentioned above, other related ones are added to clearly and comprehensively benchmark our work. For MVGRL, we also compared its original PPR augmentation. For the node classification task, we use GCA [50], GMI [28], DGI [36], CCA-SSG [46] and SpCo [21] as baselines. For the graph classification task, we use RGCL [18] and GraphCL [43] as baselines. Detailed experimental configurations are in Appendix A.
@@ 6.2 Experimental results
We present the prediction accuracy of the node classification and graph classification tasks in Table 2 and Table 3, respectively. Our comprehensive analysis reveals distinct patterns in the effectiveness of different augmentation strategies across these two task types. For node classification, DropEdge consistently achieves the best performance across multiple datasets and CG-SSL frameworks, demonstrating superior robustness and consistency. While AddEdge also achieves competitive accuracy, DropEdge stands out in this area. In graph classification, AddEdge frequently achieves the best performance across multiple datasets and CG-SSL frameworks, showing superior and more consistent results. The effectiveness of AddEdge in graph classification may be attributed to ’beneficial oversmoothing’ as proposed by [33]. In graph-level tasks, the convergence of node features to a common representation aligned with the global output can be advantageous. By potentially increasing graph density and the proportion of positively curved edges [25], AddEdge might facilitate this beneficial effect in graph classification tasks. Notably, all the results from SPAN as well as GASSER and SpCo generally underperform relative to both DropEdge and AddEdge  while also encountering scalability issues on larger datasets and suffering from a high overhead of training time.
Table 2: Node classification. Results of baselines with $\dagger$’ are adopted directly from previous works. MVGRL+PPR is the original setting of MVGRL. The best results in each cell are highlighted in grey. The best results overall are highlighted with bold and underline. Metric is accuracy (%).
@@
Table 3: Graph classification. Results of baselines with $\dagger$’ are adopted directly from previous works. MVGRL+PPR is the original setting of MVGRL. The best results in each cell are highlighted in grey. The best results overall are highlighted with bold and underline. Metric is accuracy (%).
@@
6.3 Ablation Study
To validate our findings, we conducted a series of ablation experiments on two exemplar datasets, Cora and MUTAG, representing node- and graph-level tasks, respectively. These ablation studies are crucial to rule out potential confounding variables, such as model architectures and hyperparameters, ensuring that our conclusions about the performance of CG-SSL are robust and comprehensive.
Number of Layers of GCN Encoder. To assess the impact of model depth, we conducted both node-level and graph-level experiments using varying numbers of GCN encoder layers. This analysis is to rule out the possibility that model depth, rather than augmentation strategies, influences the claim. As expected, the results, detailed in Appendix E.1, show that deeper encoders generally lead to worse performance. This suggests that excessive model complexity may introduce noise or overfitting, diminishing the benefits of spectral information. Therefore, our conclusion still holds tightly.
Type of GNN Encoder. While we initially selected GCN to align with the common protocols in previous studies for a fair comparison, we also explored other GNN architectures to ensure our findings are not specific to GCN alone. To further validate our results, we conducted additional experiments using GAT [37] for both node- and graph-level tasks, as well as GPS [30] for the graph-level task. As reported in Appendix E.2, the performance trends observed with GAT and GPS are consistent with those obtained using GCN. This consistency across different encoder types further supports our conclusion that simple edge perturbation strategies are sufficient, and that spectral augmentation does not significantly enhance performance, regardless of the type of GNN encoder applied.
@@ 7 The insignificance of Spectral Cues
Given the superior empirical performance of edge perturbations mentioned in Sec. 6, one may still argue whether it is a result of some spectral cues or not, as all the analyses mentioned are not direct evidence of the insignificance of the spectral information in the study. To clarify this, we have three questions to answer, (1) Can GNN encoders learn spectral information from augmented graphs produced edge perturbations? (2) Are spectrum in spectral augmentation necessary? (3) Is spectral information statistically a significant factor in the performance of edge perturbation? Given the questions, we conduct a series of experimental studies to answer them respectively in Sec. 7.1, 7.2 and Appendix E.3.
7.1 Degeneration of the spectrum after Edge Perturbation (EP)
@@
Here we want to conduct studies to answer the question of whether the GNN encoders applied can learn spectral information from the augmented graph views produced by EP. Therefore, we collect the spectrum of all augmented graphs ever produced along the way of the contrastive learning process of the best framework with the optimal parameter we have in this study, i.e., G-BT + EP with best drop rate $p$ or add rate $q$, and calculate the average one for each representative dataset in this study for both node- and graph-level tasks. We find that though the average spectrum of those original graphs is strikingly different, that of augmented graphs is quite similar for node- and graph-level tasks, respectively. This indicates a certain degree of degeneration of the spectra as they are no longer easy to separate after EP. Therefore, GNN encoders can hardly learn spectral information and properties between different original graphs from those augmented graph views. Note that, though we have defined some context of frameworks, this result is generally only dependent on the augmentation methods. We will elaborate on both the node-level and graph-level results in this section.
(a) Average spectra of original graphs
(b) Average spectra of augmented graphs
(c) The Spectrum of MUTAG
(d) The Spectrum of PROTEINS
@@
Node-Level Analysis. Here, we visualize the distributions of the average spectrum of graphs at the node level using histograms. The spectral distribution for each graph is represented by a sorted vector of its eigenvalues. When referring to the average spectrum, we mean the average over the eigenvalue vectors of each augmented graph. We plot the histograms of different spectra, normalized to show the probability density. Note that eigenvalues are constrained within the range [0, 2], as we adopted the commonly used symmetrical normalization. We analyze the spectral distributions of three node classification datasets: Cora, CiteSeer, and Computers. We compare the average spectral properties of both original and augmented graphs. The augmentation method used is DropEdge, applied with optimal parameters identified for the G-BT method. The results of the visualization are presented in Fig. 3. By comparing the spectrum distributions of original graphs for the datasets in Fig. 3(a), we can easily distinguish the spectra of the three datasets. This contrasts with the highly overlapped average spectra of all the datasets, indicating the degeneration mentioned. To support this claim, we also present the comparison of the spectra of original and augmented graphs on all three datasets in Fig. 3(c), 3(d), and 3(e), respectively, to show the obvious changes after the edge perturbations.
Graph-Level analysis. For graph-level analysis, we basically follow the settings mentioned above in node-level one. The only difference from the node-level task is that we have multiple original graphs with various numbers of nodes, leading to the inconsistent dimensions of the vector of the eigenvalues. Therefore, to provide a more detailed comparison of spectral properties at the graph level, we employ Kernel Density Estimation (KDE) [27] to interpolate and smooth the distributions of eigenvalues. We compare two groups of graph spectra. Each group’s spectra are processed to compute their KDEs, and the mean and standard deviation of these KDEs are calculated.
Like the results in node-level analysis, in Fig. 2(a) and 2(b), we witness the obvious difference between the average spectra of original graphs while the significant overlap between those of augmented graphs, especially if pay attention to the overlapping of the area created by the standard deviation of KDEs. Again, this contrast is not trivial because of the striking mismatch between the average spectra of original and augmented graphs in both datasets, as presented in Fig. 2(c) and 2(d).
@@ 7.2 Spectral Perturbation
Spectral perturbation on spectral augmentation baselines. SPAN, being a state-of-the-art spectral augmentation algorithm, demonstrated the correlation between graph spectra and model performance through designated perturbation on spectral priors. However, the effectiveness of simple edge perturbation motivated us to further investigate whether such a relationship is causational.
Specifically, for each pair of SPAN augmented graphs $\mathcal{G}^{1},\mathcal{G}^{2}$, we further augment them into $\mathcal{G}^{1}_{SPA},\mathcal{G}^{2}_{SPA}$ with our proposed SPA augmentor. The SPA-augmented training is performed under the same setup as SPAN, with graphs being SPA-augmented graphs $\mathcal{G}_{SPA}$. Results in Fig 4 show that the effectiveness of graph augmentation can be preserved and, in some cases improved, even if the spectral information is destroyed.
@@
$s_{SPAN}\sim p_{\theta}\left(\bm{S}_{SPAN}\left.\right|\bm{\mathcal{G}}_{0}\right),\qquad\mathcal{G}_{SPAN}\sim p_{\phi}\left(\bm{\mathcal{G}}_{SPAN}\left.\right|\bm{S}_{SPAN}\right)$ (6)
@@
$d_{SPA}\rightarrow\infty\implies p_{\theta}\left(s_{SPA}\left.\right|\bm{\mathcal{G}}_{0}\right)\rightarrow 0\implies p_{\theta,\phi}\left(\mathcal{G}_{SPA}\left.\right|\bm{\mathcal{G}}_{0}\right)\rightarrow 0$ (7)
Although the constrained generative process in Eq. 6 does indicate some extent of causality between spectral distribution $\bm{S}$ and the spectral-augmented graph distribution $\bm{\mathcal{G}}_{SPAN}$, our experiment challenges a more essential and fundamental aspect of such reasoning: such causality exists upon pre-defined generative processes, which does not intrinsically exist in the graph distributions. Even worse, such constrained generative process is incapable of modeling the full distribution of $\bm{\mathcal{G}}_{SPAN}$ itself. In our experiment setup, all $\mathcal{G}_{SPA}$ serve as strong counter examples.
@@ 8 Conclusion
In this study, we investigate the effectiveness of spectral augmentation in contrast-based graph self-supervised learning (CG-SSL) frameworks to answer the question: Are spectral augmentations necessary in CG-SSL? Our findings indicate that spectral augmentation does not significantly enhance learning efficacy. Instead, simpler edge perturbation techniques, such as random edge dropping for node-level tasks and random edge adding for graph-level tasks, not only compete well but often outperform spectral augmentations. To be specific, we demonstrate that the benefits of spectral augmentation diminish with shallower networks, and edge perturbations yield superior performance in both node- and graph-level classification tasks. Also, GNN encoders struggle to learn spectral information from augmented graphs, and perturbing edges to alter spectral characteristics does not degrade model performance. Furthermore, our theoretical analysis (Theorem  1) reveals that the InfoNCE loss bounds the mutual information achievable by augmentations, highlighting that the relatively limited direct contribution of spectral augmentations compared to simpler edge perturbations, especially when in shallow GNNs. These results challenge the current emphasis on spectral augmentation, advocating for more straightforward and effective edge perturbation techniques in CG-SSL, potentially refining the understanding and implementation of graph self-supervised learning.
Appendix A Dataset and training configuration